How do you put systems of equations into slope-intercept form?
Specifically:
-x + y = 3
-x + y = -2

Answer by Alan3354(69443)

How do you put systems of equations into slope-intercept form?
Specifically:
-x + y = 3
-x + y = -2
=========
You put each eqn into slope-int form by solving for y.